OpenStudy (anonymous):

Julie is cycling at a speed of 3.4 meters/second. If the combined mass of the bicycle and Julie is 30 kilograms, what is the kinetic energy?

OpenStudy (btaylor):

KE = 1/2 x m x v^2 KE = 1/2(30)(3.4)^2 KE = 173.4 Joules (J)